# Can LLMs’ Tuning Methods Work in Medical Multimodal Domain?

### v1

@@ Abstract
While large language models (LLMs) excel in world knowledge understanding, adapting them to specific subfields requires precise adjustments. Due to the model’s vast scale, traditional global fine-tuning methods for large models can be computationally expensive and impact generalization. To address this challenge, a range of innovative Parameters-Efficient Fine-Tuning (PEFT) methods have emerged and achieved remarkable success in both LLMs and Large Vision-Language Models (LVLMs). In the medical domain, fine-tuning a medical Vision-Language Pretrained (VLP) model is essential for adapting it to specific tasks. Can the fine-tuning methods for large models be transferred to the medical field to enhance transfer learning efficiency? In this paper, we delve into the fine-tuning methods of LLMs and conduct extensive experiments to investigate the impact of fine-tuning methods for large models on existing multimodal models in the medical domain from the training data level and the model structure level. We show the different impacts of fine-tuning methods for large models on medical VLMs and develop the most efficient ways to fine-tune medical VLP models. We hope this research can guide medical domain researchers in optimizing VLMs’ training costs, fostering the broader application of VLMs in healthcare fields. Code and dataset will be released upon acceptance.
@@ 1 Introduction
Some works[17] have attempted to apply PEFT methods to large-scale vision-language models (LVLMs). Compared to NLP tasks, visual language tasks introduce visual inputs, leading to more diverse content and requiring more challenging fine-tuning. These endeavours have demonstrated that PEFT methods successful in LLMs can enhance the few-shot and zero-shot capabilities of LVLMs or achieve comparable results to global fine-tuning approaches.
@@
To empower researchers in the medical domain with limited computational resources to effectively fine-tune multi-modal models for practical applications, we embark on experimental research to investigate the applicability of the LLMs’ tuning methods in the realm of medical multimodal learning. In this paper, we design a Modularized medIcal vision-Language fine-tuning modEl (MILE) that builds upon a medical VLP model and incorporates various PEFT modules through modular design. Specifically, from the model structure perspective, we conduct a systematic investigation of the PEFT methods in the LLMs and develop the corresponding modules which can be integrated into a generative vision-language baseline model. From the training data perspective, we propose an instruction-format medical multi-modal dataset for applying instruction-tuning on different MILE variants. We conduct in-depth ablation studies on those LLMs’ tuning methods and validate them on two radiographic image benchmarks. We believe that these empirical analyses will catalyze the development of fine-tuned multimodal models in the medical field.
Our main contributions are as follows: (i) We systematically explored how trainable parameters in different medical VLM modules affect overall performance, revealing strategies for achieving competitive results akin to global fine-tuning. (ii) Through extensive experiments, we conducted a novel comparison of the PEFT methods tailored for small-scale medical VLM based on a baseline model, offering insights distinct from large-scale models. (iii) We conducted a thorough analysis of the impact of instruction-tuning on fine-tuning basic VLP models. Our investigation revealed both positive and negative effects of instruction-tuning, offering a nuanced understanding of its implications for the fine-tuning process of the small-scale VLP models.
@@ 2 Related Work
PEFT Techniques of LLMs: Fine-tuning large pretrained language models (PLMs) is resource-intensive, often requiring substantial computational resources and training data. To address this challenge, PEFT techniques have emerged, aiming to enhance PLMs’ performance on specific tasks with minimal changes to model parameters. Various methods have been developed in this regard. Adapter Tuning[7], Dora[18], LoRA[8] etc.[16, 19] add small components to PLMs or the input embedding[10][1] to realize PEFT. Besides methods that add small components, some PEFT techniques focus on data manipulation to minimize or eliminate changes to the original model weights. OpenAI[1] and Google[23] have independently introduced instruction-tuning methods that modify original data into instruction pairs for fine-tuning models, achieving better generative results compared to multi-task training. While these PEFT methods have been successfully applied to LLMs, their impact on small-scale VLMs remains underexplored, especially in the medical domain.
Medical Vision-Language Models: Benefiting from learning-based methods [30, 24, 25, 28, 27, 31, 26, 29], recent advancements in the pretraining-finetuning paradigm have led to the emergence of medical VLMs[3, 2, 13] based on VLP models. However, their pretraining and fine-tuning require data scales of more than 100,000 image-text pairs and the number of trainable parameters for global fine-tuning is not much different from that of PEFT in LVLMs. Therefore, under the dual factors of large-scale training data and high training parameters, small-scale VLMs’ training costs remain unaffordable for many researchers. Thus, in this work, we systematically review LLMs’ tuning methods and discuss their applicability to medical VLMs.
@@ 3.1 Architecture of MILE
Baseline model: Most VLMs architecture are based on CLIP[21] or BLIP[12]. In this paper, we use MISS[2], a generative multimodal medical VLM as our baseline model, the architecture has been shown in Figure 1a. MISS has an image encoder and a JTM encoder, the former for image feature extraction and the latter for text feature extraction and multimodal feature interaction. A text decoder is appended after the JTM decoder for causal reasoning and text generation. The image encoder of the baseline model is a ViT-base[6] model; The JTM encoder is designed based on Bert with 12 transformer-based[22] layers where a cross-attention layer is inserted between the bi-self attention layer and the feed-forward layer; the architecture of the text decoder is similar to the JTM encoder and the bi-self attention layer is replaced by a causal attention layer.
The Unified Model: To validate the effectiveness of the above PEFT methods on small-scale medical VLMs, we construct MILE and equipped it with 4 modules of commonly used PEFT methods[8, 14, 16, 19], obtaining four variants: MILE-LoRA, MILE-Prefix, MILE-IA3, and MILE-PTv2.
Figure 1: Architecture of MILE with different PEFT modules(a) and PEFT details in a Transformer layer(b)
@@
$\bullet$ MILE-$(IA)^{3}$: For $(IA)^{3}$-Tuning(IA3)[16], learnable vectors which rescale the $key$, $value$ and the inner activations are respectively injected into all the attention layers and feed-forward layers in the image encoder, JTM encoder or text decoder of the baseline model during tuning, as shown in Figure 1b.
@@
$\bullet$ MILE-PTv2: As shown in Figure 1b, prompt tokens are selectively attached before the input of the JTM encoder or text decoder. Different from prefix-tuning, all prefixes in the input of each attention layer are derived from a trainable matrix when we apply P-Tuning V2[19].
@@ 4 Experiment Results and Analysis
We initially trained a series of MILE equipped with PEFT modules using the original data. Tables 1 to 4 present the accuracy(ACC(%)) of four MILE variants training with the origin data on the Slake benchmark. ’F’ denotes the freezing of all parameters within a given module, ’T’ signifies that all parameters are trainable, and the acronyms ’LoRA’, ’Prefix’, ’IA3’, and ’PTV2’ identify the specific PEFT methods applied.
Table 1: Results of MILE-LoRA(origin data)
ViT JTM Dec Rank #Params Opened Closed Gobal
LoRA 4 0.163% 3.57 50.70 20.34
F LoRA LoRA 8 0.325% 3.57 50.70 20.34
4 0.327% 48.65 50.70 49.34
LoRA LoRA LoRA 8 0.652% 48.93 50.70 49.57
LoRA 4 38.022% 47.76 70.70 55.53
F T LoRA 8 38.072% 50.21 70.99 57.18
LoRA 4 24.009% 68.14 50.70 62.29
T LoRA LoRA 8 24.133% 68.28 50.70 62.38
4 61.887% 78.52 79.44 78.83
T T LoRA 8 61.919% 78.66 80.56 79.30
@@
Table 2: Results of MILE-Prefix. ViT JTM Dec #Params Opened Closed Global F F Prefix 3.926% 0 50.7 17.3 F Prefix Prefix 7.556% 0 50.7 17.3 T Prefix Prefix 29.636% 41.50 32.95 38.61 T T Prefix 63.354% 76.82 82.25 78.65 Table 3: Results of MILE-IA3. ViT JTM Dec #Params Opened Closed Global F IA3 IA3 0.051% 0 1.69 0.57 IA3 IA3 IA3 0.061% 0 50.70 16.98 T IA3 IA3 23.924% 12.77 28.17 17.92 F T IA3 37.987% 46.24 50.70 47.74 T T IA3 61.866% 72.20 47.04 63.77 Table 4: Results of MILE-PTV2. ViT JTM Dec #Params Opened Closed Global F PTV2 PTV2 0.102% 0 0 0 F F PTV2 0.051% 7.10 0 4.72 T PTV2 PTV2 23.963% 13.62 29.30 18.87 T T PTV2 61.876% 74.18 49.86 66.04
[Uncaptioned image] Figure 2: ACC of MILE trained with different data.
@@ 4.1 Performance Differences Among Different PEFT Methods
Furthermore, we conducted an ablation study on the rank of the LoRA unit. As shown in Table 1, within the same tuning paradigm, a doubling of the number of parameters in the LoRA matrix brings about a minor improvement in model performance. A LoRA rank of 8(ViT freezed) and LoRA-tuned(rank=4) visual encoders, adjusting a similar parameter fraction (about 0.32%), differed by 45% in open-ended questions’ accuracy.
@@
Figure 3: Loss of MILE variants under the training with origin data(a), instruction-format data(b) and a instruction-tuning MILE after fine-tuning by origin data(c)
@@ 4.3 Is Instruction-Tuning Ineffective for Basic VLMs?
As shown in Figure 3, training with instruction-based data showed lower losses, hinting at an improved task format comprehension but potentially oversimplifying the difficulty of training tasks, thus diminishing real-world inferencing efficacy.
However, given that instruction-format data can enhance the model’s understanding of the target task and result in lower training losses, could it potentially improve a global fine-tuned model that has already converged on the original data? Figure 3 demonstrates the loss reduction when a converged model is further trained on instruction-format data, the loss of the converged model continues to decrease and finally reaches about 4.33 after global fine-tuning. As shown in Table 5, MILE ultimately achieves 86.7% open-ended ACC, 81.42% closed-ended ACC, and 83.02% overall ACC, surpassing the baseline model and demonstrating state-of-the-art performance among generative VLMs.
Table 5: Comparsion with other medical VLMs which have different tuning paradigms, $\clubsuit$’ means global fine-tuning with ordinary data(no instruction-format) and $\spadesuit$’ means instruction-tuning with all parameters updating
@@ 1 Templates of instruction-format data
Fine-tuning with instructions has demonstrated promising results in both language models and LVLMs. However, there is still a lack of research on the impact of fine-tuning on fundamental VLMs, especially in the medical domain. We constructed an instruction-format medical image-text dataset to investigate their impact on fundamental VLMs. Drawing inspiration from Instruction-BLIP, We transformed the QA pairs into an instruction fine-tuning format by constructing instruction templates, as shown in Figures 4 to 5. ’Closed’ templates are designed for closed-ended questions and ’Opened’ templates are suitable for open-ended questions. Each QA pair is randomly embedded with one of these templates during training. To generate answer options, we classified the question attributes based on modality, plane, shape, size, organ, location, and pathology. This allowed us to create pools of candidate answers for different question attributes. When creating opened instructions, incorrect answers from the same attribute are randomly selected and embedded into the option together with the ground-truth answer.
Figure 4: Instruction-format Data Templates
Figure 5: Instruction-format Data Examples
@@ 2 Details of Experiments
More Detailed Results of MILE Variants trained with Instruction-format Data: As mentioned in Chapter 4, we also trained a series of MILE equipped with PEFT modules using the instruction-format data. Tables 6 to 7 present the accuracy(ACC) of four MILE variants on the benchmark.
Table 6: Results of MILE-LoRA(instruction-format data)
ViT JTM Dec Rank #Params Opened Closed Gobal
LoRA 4 0.163% 0 50.7 16.98
F LoRA LoRA 8 0.325% 0 50.7 16.98
4 0.327% 28.09 29.01 28.40
LoRA LoRA LoRA 8 0.652% 48.93 24.82 31.51
LoRA 4 38.022% 21.79 35.77 26.49
F T LoRA 8 38.072% 27.35 39.44 31.32
LoRA 4 24.009% 39.57 8.73 29.25
T LoRA LoRA 8 24.133% 41.42 23.10 35.28
4 61.887% 61.28 36.34 52.92
T T LoRA 8 61.919% 63.54 45.35 57.45
Table 7: Results of MILE-Prefix & IA3 & PTV2(instruction-format data)
ViT JTM Dec #Params Opened Closed Gobal
F IA3 IA3 0.051% 0 50.70 16.98
IA3 IA3 IA3 0.061% 0 50.70 16.98
T IA3 IA3 23.924% 12.77 27.04 17.92
F T IA3 37.987% 8.37 27.04 14.62
T T IA3 61.866% 50.21 49.86 50.09
F F Prefix 3.926% 0 50.70 17.30
F Prefix Prefix 7.556% 0 50.70 17.30
T Prefix Prefix 29.636% 7.23 22.38 12.64
T T Prefix 63.354% 68.65 32.39 56.51
F F PTV2 0.051% 7.10 0 4.72
F PTV2 PTV2 0.102% 0 0 0
T PTV2 PTV2 23.963% 6.10 23.38 11.89
T T PTV2 61.876% 3.12 30.99 12.43
Implement Details: Here, we will present the experimental details of our model. All the training was conducted on a single NVIDIA RTX8000-48GB GPU. We used the Adamw optimizer with cosine learning rate decay, an initial learning rate of 2e-5, a weight decay of 0.05, and a minimum learning rate of 0. The input image size for our model was 480x480 pixels, trained for 130 epochs, with a batch size of 20. For our baseline model, the Vit-based visual encoder consists of 12 layers of transformer, and both the JTM encoder and text decoder contain 12 layers of transformer-based layers.
Why choose this baseline model: We choose MISS as our baseline model and develope MILE based on it because MISS is a generative VLM and it choose ALBEF, a mainstream framework as its architecture. MISS unifies the Text encoder and the Multimodal encoder by a JTM encoder so we can more easily fine-tune it.

### v2

@@ Abstract
While Large Language Models (LLMs) excel in world knowledge understanding, adapting them to specific subfields requires precise adjustments. Due to the model’s vast scale, traditional global fine-tuning methods for large models can be computationally expensive and impact generalization. To address this challenge, a range of innovative Parameters-Efficient Fine-Tuning (PEFT) methods have emerged and achieved remarkable success in both LLMs and Large Vision-Language Models (LVLMs). In the medical domain, fine-tuning a medical Vision-Language Pretrained (VLP) model is essential for adapting it to specific tasks. Can the fine-tuning methods for large models be transferred to the medical field to enhance transfer learning efficiency? In this paper, we delve into the fine-tuning methods of LLMs and conduct extensive experiments to investigate the impact of fine-tuning methods for large models on the existing multimodal model in the medical domain from the training data level and the model structure level. We show the different impacts of fine-tuning methods for large models on medical VLMs and develop the most efficient ways to fine-tune medical VLP models. We hope this research can guide medical domain researchers in optimizing VLMs’ training costs, fostering the broader application of VLMs in healthcare fields. The code and dataset have been released at https://github.com/TIMMY-CHAN/MILE.
@@ 1 Introduction
Some works [18, 2] have attempted to apply PEFT methods to Large Vision-Language Models (LVLMs). Compared to Natural Language Processing (NLP) tasks, visual language tasks introduce visual inputs, leading to more diverse content and requiring more challenging fine-tuning. These endeavours have demonstrated that PEFT methods successful in LLMs can enhance the few-shot and zero-shot capabilities of LVLMs or achieve comparable results to global fine-tuning approaches.
@@
To empower researchers in the medical domain with limited computational resources to effectively fine-tune multi-modal models for practical applications, we embark on experimental research to investigate the applicability of the LLMs’ tuning methods in the realm of medical multimodal (vision-language) learning. In this paper, we design a Modularized medIcal vision-Language fine-tuning modEl (MILE) that builds upon a medical Vision-Language Pretrained (VLP) model and incorporates various PEFT modules through modular design. Specifically, from the model structure perspective, we conduct a systematic investigation of the PEFT methods in the LLMs and develop the corresponding modules which can be integrated into a generative vision-language baseline model. From the training data perspective, we propose an instruction-format medical multi-modal dataset for applying instruction-tuning on different MILE variants. We conduct in-depth ablation studies on those LLMs’ tuning methods and validate them on two radiographic image benchmarks. We believe that these empirical analyses will catalyze the development of fine-tuned medical multimodal models.
Our main contributions are as follows: (i) We systematically explored how trainable parameters in different medical VLM modules affect overall performance, revealing strategies for achieving competitive results akin to global fine-tuning. (ii) Through extensive experiments, we conducted a novel comparison of the PEFT methods tailored for small-scale medical VLM based on a baseline model, offering insights distinct from large-scale models. (iii) We conducted a thorough analysis of the impact of instruction-tuning on fine-tuning basic VLP models and released an instruction-format medical image-text dataset. Our investigation revealed both positive and negative effects of instruction-tuning, offering a nuanced understanding of its implications for the fine-tuning process of the small-scale VLP models.
@@ 2 Related Work
PEFT Techniques of LLMs: Fine-tuning large pretrained language models (PLMs) is resource-intensive, often requiring substantial computational resources and training data. To address this challenge, PEFT techniques have emerged, aiming to enhance PLMs’ performance on specific tasks with minimal changes to model parameters. Various methods have been developed in this regard. Adapter Tuning [8], Dora [19], LoRA [9] etc. [17, 20] add small components to PLMs or the input embedding [11, 1] to realize PEFT. Besides methods that add small components, some PEFT techniques focus on data manipulation to minimize or eliminate changes to the original model weights. OpenAI [1] and Google [24] have independently introduced instruction-tuning methods that modify original data into instruction pairs for fine-tuning models, achieving better generative results compared to multi-task training. While these PEFT methods have been successfully applied to LLMs, their impact on small-scale VLMs remains underexplored, especially in the medical domain.
Medical Vision-Language Models: Recent advancements in the pretraining-finetuning paradigm have led to the emergence of medical VLMs [4, 3, 14] based on VLP models. However, their pretraining and fine-tuning require data scales of more than 100,000 image-text pairs and the number of trainable parameters for global fine-tuning is not much different from that of PEFT in LVLMs. Therefore, under the dual factors of large-scale training data and high training parameters, small-scale VLMs’ training costs remain unaffordable for many researchers. Thus, in this work, we systematically review LLMs’ tuning methods and discuss their applicability to medical VLMs.
@@ 3.1 Architecture of MILE
Baseline model: Most VLMs architecture are based on CLIP [22] or BLIP [13]. In this paper, we use MISS[3], a generative multimodal medical VLM as our baseline model, the architecture has been shown in Figure 1a. MISS has an image encoder and a Joint Text-Multimodal (JTM) encoder, the former for image feature extraction and the latter for text feature extraction and multimodal feature interaction. A text decoder is appended after the JTM decoder for causal reasoning and text generation. The image encoder of the baseline model is a ViT-base[7] model; The JTM encoder is designed based on Bert with 12 transformer-based[23] layers where a cross-attention layer is inserted between the bi-self attention layer and the feed-forward layer; the architecture of the text decoder is similar to the JTM encoder and the bi-self attention layer is replaced by a causal attention layer.
The Unified Model: To validate the effectiveness of the above PEFT methods on small-scale medical VLMs, we construct MILE and equip it with 4 modules of commonly used PEFT methods [9, 15, 17, 20], obtaining four variants: MILE-LoRA, MILE-Prefix, MILE-IA3, and MILE-PTv2.
Figure 1: Architecture of MILE with different PEFT modules (a) and PEFT details in a Transformer layer (b).
@@
$\bullet$ MILE-$(IA)^{3}$: For $(IA)^{3}$-Tuning (IA3) [17], learnable vectors $l_{k}$, $l_{v}$ and $l_{ff}$ which rescale the $key$, $value$ and the inner activations are respectively injected into all the attention layers and feed-forward layers in the image encoder, JTM encoder or text decoder of the baseline model during tuning, as shown in Figure 1b.
@@
$\bullet$ MILE-PTv2: As shown in Figure 1a, prompt tokens are selectively attached before the input of the JTM encoder or text decoder. Different from prefix-tuning, all prefixes in the input of each attention layer are derived from a trainable matrix when we apply P-Tuning v2[20].
@@ 4 Experiment Results and Analysis
We initially trained a series of MILE equipped with PEFT modules using the original data. Tables 1 to 4 present the accuracy (ACC(%)) of four MILE variants training with the origin data on the Slake benchmark. ’F’ denotes the freezing of all parameters within a given module, ’T’ signifies that all parameters are trainable, and the acronyms ’LoRA’, ’Prefix’, ’IA3’, and ’PTV2’ identify the specific PEFT methods applied. ’Memory’ represents the GPU memory (GB) required for training. And ’#Params’ indicates the weight of trainable parameters over all parameters.
Table 1: Results of MILE-LoRA(origin data).
ViT JTM Dec Rank #Params Memory Opened Closed Gobal
LoRA 4 0.163% 5.19 3.57 50.70 20.34
F LoRA LoRA 8 0.325% 5.21 3.57 50.70 20.34
4 0.327% 26.63 48.65 50.70 49.34
LoRA LoRA LoRA 8 0.652% 26.75 48.93 50.70 49.57
LoRA 4 38.022% 7.26 47.76 70.70 55.53
F T LoRA 8 38.072% 7.45 50.21 70.99 57.18
LoRA 4 24.009% 26.96 68.14 50.70 62.29
T LoRA LoRA 8 24.133% 27.29 68.28 50.70 62.38
4 61.887% 27.60 78.52 79.44 78.83
T T LoRA 8 61.919% 28.11 78.66 80.56 79.30
@@
Table 2: Results of MILE-Prefix.
ViT JTM Dec #Params Memory Opened Closed Global F F Prefix 3.926% 4.62 0 50.7 17.3 F Prefix Prefix 7.556% 4.67 0 50.7 17.3 T Prefix Prefix 29.636% 26.41 41.50 32.95 38.61 T T Prefix 63.354% 27.97 76.82 82.25 78.65
Table 3: Results of MILE-IA3.
ViT JTM Dec #Params Memory Opened Closed Global F IA3 IA3 0.051% 6.35 0 1.69 0.57 IA3 IA3 IA3 0.061% 23.01 0 50.70 16.98 T IA3 IA3 23.924% 26.83 12.77 28.17 17.92 F T IA3 37.987% 7.52 46.24 50.70 47.74 T T IA3 61.866% 27.90 72.20 47.04 63.77
Table 4: Results of MILE-PTV2.
ViT JTM Dec #Params Memory Opened Closed Global F PTV2 PTV2 0.102% 4.52 0 0 0 F F PTV2 0.051% 4.57 7.10 0 4.72 T PTV2 PTV2 23.963% 25.41 13.62 29.30 18.87 T T PTV2 61.876% 27.46 74.18 49.86 66.04
Figure 2: ACC of MILE trained with different data.
@@ 4.1 Performance Differences Among Different PEFT Methods
Furthermore, we conducted an ablation study on the rank of the LoRA unit. As shown in Table 1, within the same tuning paradigm, a doubling of the number of parameters in the LoRA matrix brings about a minor improvement in model performance. A LoRA rank of 8 (ViT froze) and LoRA-tuned (rank=4) visual encoders, adjusting a similar parameter fraction (about 0.32%), differed by 45% in open-ended questions’ accuracy.
@@
Figure 3: Loss of MILE variants under the training with origin data (a), instruction-format data (b) and an instruction-tuning MILE after fine-tuning by origin data (c).
@@ 4.3 Is Instruction-Tuning Ineffective for Basic VLMs?
As shown in Figure 3a&b, global fine-tuning with instruction-based data showed lower losses, hinting at an improved task format comprehension but potentially oversimplifying the difficulty of training tasks, thus diminishing real-world inferencing efficacy.
However, given that instruction-format data can enhance the model’s understanding of the target task and result in lower training losses, could it potentially improve a global fine-tuned model that has already converged on the original data? Figure 3c demonstrates the loss reduction when a converged model is further trained on instruction-format data, the loss of the converged model continues to decrease and finally reaches about 4.33 after global fine-tuning. As shown in Table 5, MILE ultimately achieves 86.7% open-ended ACC, 81.42% closed-ended ACC, and 83.02% overall ACC, surpassing the baseline model and demonstrating state-of-the-art performance among generative VLMs.
Table 5: Comparsion with other medical VLMs which have different tuning paradigms, $\clubsuit$’ means global fine-tuning with ordinary data (no instruction-format) and $\spadesuit$’ means instruction-tuning with all parameters updating.
@@ 5 Conclusion
6 Acknowledgement
This work is supported in part by the National Key R&D Program of China (2021ZD0113502).
@@ 1 Templates of instruction-format data
We constructed an instruction-format medical image-text dataset to investigate their impact on fundamental VLMs. In detail, we transformed the QA pairs into an instruction fine-tuning format by constructing instruction templates, as shown in Figures 4 to 5. ’Closed’ and ’Opened’ templates are designed for closed-ended and open-ended questions respectively. Each QA pair is randomly embedded with one of these templates during training. To generate answer options, we classified the question attributes based on modality, plane, shape, size, organ, location, and pathology. Consequently, we created pools of candidate answers for different question attributes. Incorrect answers from the same attribute are randomly selected and embedded into the option together with the ground-truth answer.
Figure 4: Instruction-format Data Templates.
Figure 5: Instruction-format Data Examples.
@@ 2 Details of Experiments
More Detailed Results of MILE Variants: As mentioned in Section 4 of our paper, we trained a series of MILE variants using the instruction-format data, and Tables 6 to 7 present the related results of four MILE variants on the Slake dataset. To more fully demonstrate the impact of different PEFT methods on the basic visual language model, we have also verified the PEFT method’s impacts on BiomedGPT-Tiny: only the decoder is fine-tuned with PEFT methods, and the rest is full parameters fine-tuned. Table 8 presents the ACC on the benchmark, which shows a similar trend as MILE.
Implement Details: Here, we will present the experimental details of our model. All the training was conducted on a single NVIDIA RTX8000-48GB GPU. We used the Adamw optimizer with cosine learning rate decay, an initial learning rate of 2e-5, a weight decay of 0.05, and a minimum learning rate of 0. The input image size for our model was 480$\times$480 pixels, trained for 130 epochs, with a batch size of 20. For our baseline model, the Vit-based visual encoder consists of 12 layers of transformer, and both the JTM encoder and text decoder contain 12 layers of transformer-based layers.
Why choose this baseline model: We chose MISS as our baseline model and developed MILE based on it because MISS is a small-scale generative VLM and it has similar architecture to current LVLMs such as LLaVA, BLIP2, and et.al. MISS unifies the Text encoder and the Multimodal encoder by a JTM encoder so we can more easily fine-tune it.
Table 6: Results of MILE-LoRA (instruction-format data).
ViT JTM Dec Rank #Params Memory Opened Closed Gobal
LoRA 4 0.163% 5.44 0 50.7 16.98
F LoRA LoRA 8 0.325% 5.53 0 50.7 16.98
4 0.327% 26.90 28.09 29.01 28.40
LoRA LoRA LoRA 8 0.652% 27.01 48.93 24.82 31.51
LoRA 4 38.022% 7.85 21.79 35.77 26.49
F T LoRA 8 38.072% 7.94 27.35 39.44 31.32
LoRA 4 24.009% 26.95 39.57 8.73 29.25
T LoRA LoRA 8 24.133% 27.62 41.42 23.10 35.28
4 61.887% 27.66 61.28 36.34 52.92
T T LoRA 8 61.919% 28.63 63.54 45.35 57.45
Table 7: Results of MILE-Prefix & IA3 & PTV2 (instruction-format data).
ViT JTM Dec #Params Memory Opened Closed Gobal
F IA3 IA3 0.051% 6.41 0 50.70 16.98
IA3 IA3 IA3 0.061% 23.47 0 50.70 16.98
T IA3 IA3 23.924% 27.42 12.77 27.04 17.92
F T IA3 37.987% 8.18 8.37 27.04 14.62
T T IA3 61.866% 28.81 50.21 49.86 50.09
F F Prefix 3.926% 4.76 0 50.70 17.30
F Prefix Prefix 7.556% 4.81 0 50.70 17.30
T Prefix Prefix 29.636% 26.56 7.23 22.38 12.64
T T Prefix 63.354% 28.14 68.65 32.39 56.51
F F PTV2 0.051% 4.66 7.10 0 4.72
F PTV2 PTV2 0.102% 4.70 0 0 0
T PTV2 PTV2 23.963% 26.03 6.10 23.38 11.89
T T PTV2 61.876% 27.83 3.12 30.99 12.43
Table 8: Results on BiomedGPT-Tiny (origin data).
Method #Params Opened Closed Global
Full Fine-tuning 100% 71.84 64.46 68.97
Decoder-LoRA 50.76% 66.82 63.48 65.52
Decoder-Prefix 51.05% 69.94 60.54 66.29
Decoder-IA3 50.49% 64.95 52.21 60.01
Decoder-PTV2 50.92% 68.07 48.78 60.57